# Magnetic properties of Cr8 and V8 molecular rings from ab initio calculations

### v1

@@ I Introduction
Molecular nanomagnets (MNM) are attracting a vast interest in the scientific community, due to their huge potential for technological applications [1], including spintronics [2, 3, 4, 5, 6], magneto-refrigeration [7, 8, 9, 10], miniaturisation of data-storage media [11], quantum sensing [12] and quantum-information [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. The reduced size of these systems and, in many cases, the availability of experimental data for comparison, make them ideal to even investigate fundamental questions such as the different behavior of integer and half-integer spin systems [24, 25, 26, 27, 28, 29, 30, 31], the influence of geometrical curvature on the spin order and magnetic interactions [32, 33], the origin of generalized spin-orbit couplings [34], the relation between spin and transport properties [35].
Here, we consider a specific class of MNM: ring-shaped molecules with eight magnetic centers (transition-metal - TM - ions with localized spins) occupied by either Cr or V, known as Cr8 and V8, respectively. Due to their modular and stable structure (TM-centered anion octahedra glued together by organic ligands) and the interchangeable TM ions, they offer a platform to control a variety of magnetic behaviors, which could be engineered for target applications. In fact, these systems have been studied for over two decades, both experimentally [36, 37, 38, 39, 40] and theoretically [41, 42, 43], with particular attention to Cr-based rings, due to their potential for quantum information technologies [44, 45, 46, 47, 48, 49, 50].
@@
In this work, we present a first computational study of V8 aimed at investigating, in comparison with the well-known Cr8, its peculiar ground state magnetic properties. Density functional theory [63, 64] calculations are used to determine the electronic ground state of both systems and compare the energy of various spin configurations. Precise values of the exchange couplings between the TM ions are obtained from DFT, upon fixing the sign (i.e., the direction) of the spin on each magnetic site, while their magnitude remains unchanged.
@@ II Cr8 and V8 molecular complexes
Unlike Cr8, the specific V8 ring we consider in this work has never been synthesized, due to parasitic oxidation processes [39]. Therefore, a direct comparison of our results with experiments is currently not feasible. To the best of our knowledge, the only two homometallic rings containing eight V magnetic centers ever synthesized and characterized experimentally are reported in Refs. [61] and [62]. However, compared to our V8, they feature a different organic scaffold around the TM centers, and, quite interestingly, report different ground-state magnetic configurations (AFM and FM, respectively). On the other hand, two heterometallic V-based rings with the same structure as our V8 (see Fig. 2) have been synthesized, with single Ni and Zn impurities (V7Ni and V7Zn, respectively). Their experimental characterization shows an AFM ground state [65, 39, 67] which has led authors to expect a similar behavior also for V8. A comparison of these systems with the ones studied here is left for future work.
@@ III.2 Exchange interactions
The magnetic couplings are evaluated from ab initio calculations, by fitting the DFT total energy of suitably chosen spin configurations with the predictions of the spin Hamiltonian. Besides widely used local and semilocal approximations to the exchange-correlation (xc) functional (such as LDA[63, 64] and GGA [68]) we use the popular DFT+U [69, 70] and the extended DFT+U+V [71, 72] approaches, necessary to improve the $d$ states localization and achieve a more accurate representation of the magnetic moments and their interactions.
@@
Here we follow a somewhat simpler approach based on the broken-symmetry DFT method (BS-DFT)  [79, 80, 81, 82, 83, 84, 85, 86, 87]. More specifically, we adopt the projected BS-DFT, a method which relies on the same assumptions of the four-states method [76, 77, 78]. The approach consists in mapping the DFT total energies of various spin configurations onto the expectation values of the Hamiltonian (Eq. 1) on states that resemble the ones achieved from DFT (i. e. single-determinants constructed from the appropriate eigenstates of $\hat{S}^{z}$ on each magnetic ion). In practice, if we indicate these states as |$\ket{\sigma_{1}^{\left(k\right)},\dots,\sigma_{8}^{\left(k\right)}}$ (where $\sigma_{i}^{\left(k\right)}$ is the eigenvalue of $\hat{S}^{z}$ on the $i^{th}$ magnetic site for the $k^{th}$ spin configuration) our approach corresponds to identifying $E_{DFT}^{(k)}$ with $\bra{\sigma_{1}^{\left(k\right)},\dots,\sigma_{8}^{\left(k\right)}}\hat{H}_{ex}\ket{\sigma_{1}^{\left(k\right)},\dots,\sigma_{8}^{\left(k\right)}}$. When a sufficiently large number of states is considered (typical pools of states include FM, AFM and some intermediate configurations), the effective exchange parameters of $\hat{H}_{ex}$ can be extracted from the best fit of the DFT total energies.
@@
After having computed the magnetic couplings, it is worth stressing that, while the structure of the molecule possesses a D4 symmetry (due to the alternate modulation of ligands groups), the magnetic centers are found to be approximately equivalent and deviations from a D8 symmetry can be safely neglected both for the modulus of the localized spins and for the interactions between them. In addition, due to the annular geometry of the systems considered here, periodic boundary conditions are assumed on the site indexes: $i+8=i$.
@@ III.3 Computational details
Density Functional Theory (DFT) calculations were performed using the publicly available, plane-wave pseudopotential package Quantum ESPRESSO (QE) [88, 89, 90]. Spin-resolved, Hubbard-corrected calculations were based on a collinear-spin formulation of DFT. Specifically, we compare results obtained from two most popular approximations for the exchange-correlation (xc) functional: the local density approximation (LDA) [63, 64] and the PBE flavor of the generalized-gradient approximation (GGA) [68]. The pseudopotentials of the TMs were obtained from the QE PSlibrary archive [91] [92] and were chosen to be Ultrasoft (US) and projected-augmented wave (PAW) for GGA and LDA, respectively. Kohn-Sham wavefunctions and charge density were expanded in plane-waves up to a kinetic-energy cutoffs of 80Ry and 640Ry, respectively, in all cases.
Due to the complexity of the real molecular systems, a simplification of the structures was necessary to make the calculations affordable. This was achieved using the widely adopted “hydrogen saturation” procedure [53, 41, 43], which consists in replacing terminal methyl (-CH3) groups with single H atoms. The procedure was sequentially applied twice, and the resulting structure can be observed in Fig. 5 (to be compared with Fig. 2).
Figure 5: Reduced molecular structure (see text for explanation).
@@
In order to improve the $d$ states localization and achieve a better representation of magnetic properties, we also used the DFT+U and DFT+U+V schemes. One-site $U$ and inter-site $V$ Hubbard parameters were computed using linear-response theory [93] within the automatic approach introduced in Refs. [94, 95, 96, 97]. To speed-up this calculation, a linear model of the rings’ local structure was used, i.e. a chain containing only two magnetic centers (instead of eight) with the same octahedral environment of the TM centers as in the whole molecule.
Figure 6: Linearized chain model of the ring molecule, used to compute the Hubbard parameters. The vertical lines delimitate the two-magnetic-sites unit cell (see text).
The chain is obtained from the simplified ring in Fig. 5 by changing the position of the ligand groups around the magnetic centers in such a way they can be arranged in a linear conformation. In the process, the structure of the local O-F octahedra around the TMs is maintained as close as possible to the original one. Test calculations of the Hubbard parameters in the ring (optimized with GGA) and in the chain yielded almost coincident results  [^1] A more exhaustive discussion about this procedure (and the importance of a comparison between curved and linear molecules) will be given in a separate publication.
In all cases the calculation of these parameters was performed self-consistently with respect to the electronic state (i.e. $U$’s and $V$’s were obtained perturbing a DFT+U(+V) ground state until input and output values converged). The obtained values are reported in Table 1, showing, for both systems, the considered xc functionals and flavours of the Hubbard correction (DFT+U and DFT+U+V), the value of the computed Hubbard parameters.
@@
It is important to remark that the results shown in Table 1 refer to the ground-state spin configuration of each system. However, test calculations have shown that the Hubbard parameters’ dependence on the arrangement of magnetic moments can be safely neglected. Table 1 shows relatively minor differences (0.1 - 0.2 eV) in the value of $U$ and $V$ between LDA and GGA, independently of the flavor of the Hubbard correction. In contrast, turning on the inter-site interaction ($V$) leads to a significant upward shift ($\approx$ 1 eV) of the on-site Hubbard $U$. At the same time the intersite $V$ exhibits surprisingly similar values for Cr and V, probably due to the close resemblance of the ligands’ structure around the TMs in the two systems. The on-site Hubbard parameters show instead a more pronounced difference between the two systems with those of Cr $\approx$ 0.6 eV higher than those of V. Finally, LDA yields higher values than GGA; the shift (non uniform and slightly more pronounced for $U$ than for $V$) is of the order of 0.1 eV.
@@ IV Results
where the additive constant $c$ fixes the energy reference. The considered spin configurations are identified by the sign of the spin on each magnetic site (the modulus being always maximum: $|\sigma_{i}|=S_{i}$) which is controlled by the proper initialization of the local magnetization. Typically, between the FM and the AFM states we consider five/six intermediate configurations (some of which are exemplified in Fig. 4) featuring various numbers of spin-flips in various positions.
A first rough estimate of the nearest-neighbor couplings can be also obtained from the direct comparison of the FM and AFM spin arrangements (see Fig. 4). The results for LDA and GGA are reported in Table 2, showing
@@
that both functionals yield quite large exchange couplings [^3] We stress that the results shown in Table 2 are rather sensitive to the molecule’s structure (i.e., can change significantly for minor readjustments of the TM centers’ coordination shell) which, in turn, is often hard to optimize. This sensitivity disappears when using DFT+U or DFT+U+V. We thus believe it to be another effect of the overdelocalization of valence $d$ electrons in LDA or GGA.. In the case of Cr8, the values strongly overestimate the experimental one ($\sim$1.46 meV, from inelastic neutron scattering (INS) data [58] and electron paramagnetic resonance (EPR) measurements [38, 36]). The mismatch is larger in LDA, which for V8 yields an unrealistic $\sim$-38 meV value. Besides these quantitative inaccuracies, it is important to notice the difference in the sign of the couplings between V8 and Cr8. In agreement with literature [57, 38, 36, 58, 43, 52, 59, 53, 54] Cr8 is found to be AFM, driven by positive nearest-neighbor exchange interactions $J^{(1)}$. In contrast, a completely different picture emerges for V8, exhibiting a negative $J^{\left(1\right)}$ that stabilizes a FM order.
Importantly, using Hubbard corrections strongly modifies the exchange interactions. In the following we present the values obtained with DFT+U and DFT+U+V (DFT is either LDA or GGA) for both systems. In this case the fit on DFT total energy was based on a larger number of spin configurations (see Fig. 4 for a selection) which also gave access to the exchange couplings between second nearest neighbors (those between further neighbors were found negligible in all cases).
@@
Table 3 collects the couplings obtained from LDA+U and GGA+U (on-site corrections only). While the couplings upon the Hubbard correction are hugely renormalized as compared to LDA and GGA, the signs of the nearest neighbor couplings are not affected, meaning that the character of the ground state in both systems remains the same: AFM for Cr8 ($J^{\left(1\right)}$ $>$ 0) and FM for V8 ($J^{\left(1\right)}$ $<$ 0). The values of $J^{\left(1\right)}$ are now of the expected order of magnitude and are much closer to the experimental value for Cr8 ($\sim$1.46 meV [58, 36, 38]). Both LDA+U and GGA+U show somewhat underestimated values.
Moreover, besides a FM $J^{\left(1\right)}$, V8 also features a sizeable second-nearest neighbor coupling $J^{\left(2\right)}$, that is positive and, in magnitude, represents a significant fraction of $J^{\left(1\right)}$. In fact, a sizeable $J^{\left(2\right)}$ is needed in order to fit the DFT total energies and make the system ”Heisenberg-like” (i.e., with couplings independent from the spin configuration). $J^{\left(2\right)}$ is also present for Cr8 but is one order of magnitude smaller than for V8. The presence of a relatively large and positive $J^{\left(2\right)}$ is an interesting feature of V8: competing (opposite sign) magnetic interactions of comparable strength might induce frustration and/or enhance the system’s susceptibility to external fields, enabling easier switching between different magnetic orders, an aspect that could be exploited in some applications.
@@
The comparison with Table 3 results clarifies immediately that the coupling $V$ (between the TMs and the anions in their first shell of neighbors) generally shifts the value of couplings upward (negative ones become less negative). This slightly improves the agreement for Cr8 with experiments compared to the DFT+U case. The shift is quantitatively more consistent for $J^{\left(1\right)}$ than $J^{\left(2\right)}$ and slightly more pronounced for LDA than GGA. Aside from these differences, however, the overall picture remains unchanged: positive (AFM) $J^{\left(1\right)}$’s and negligible $J^{\left(2\right)}$ for Cr8; negative (FM) $J^{\left(1\right)}$ and consistent, positive $J^{\left(2\right)}$ for V8.
In spite of their structural and chemical similarity (Cr and V are right next to each-other on the periodic table) the two molecules exhibit radically different ground states with an AFM order for Cr8 (mainly stabilized by $J^{\left(1\right)}$) and a FM arrangement for V8 that results from the competition between $J^{\left(1\right)}$ (FM, and numerically prevalent) and $J^{\left(2\right)}$ (AFM).

### v2

@@ I Introduction
Molecular nanomagnets (MNM) are attracting a vast interest in the scientific community, due to their huge potential for technological applications [1], including spintronics [2, 3, 4, 5, 6], magneto-refrigeration [7, 8, 9, 10], miniaturisation of data-storage media [11], quantum sensing [12] and quantum-information [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. The reduced size of these systems and, in many cases, the availability of experimental data for comparison, make them ideal to even investigate fundamental questions such as the different behavior of integer and half-integer spin systems [27, 28, 29, 30, 31, 32, 33], the influence of geometrical curvature on the spin order and magnetic interactions [34, 35], the origin of generalized spin-orbit couplings [36], the relation between spin and transport properties [37].
Here, we consider a specific class of MNM: ring-shaped molecules with eight magnetic centers consisting of transition-metal (TM) ions carrying localized spins, occupied by either Cr or V, known as Cr8 and V8, respectively. Due to their modular and stable structure composed of TM-centered anion octahedra glued together by organic ligands, and the interchangeable TM ions, they offer a platform to control a variety of magnetic behaviors, which could be engineered for target applications. In fact, these systems have been studied for over two decades, both experimentally [38, 39, 40, 41, 42] and theoretically [43, 44, 45], with particular attention to Cr-based rings, due to their potential for quantum information technologies [46, 47, 48, 49, 50, 51, 52].
@@
In this work, we present a first computational study of V8 aimed at investigating, in comparison with the well-known Cr8, its peculiar ground state magnetic properties. Density functional theory [63, 64] calculations are used to determine the electronic ground state of both systems and compare the energy of various spin configurations. Precise values of the exchange couplings between the TM ions are obtained from DFT, upon fixing the sign of the spin on each magnetic site, while their magnitude remains unchanged.
@@ II Cr8 and V8 molecular complexes
Unlike Cr8, the specific V8 ring we consider in this work has never been synthesized, due to parasitic oxidation processes [40]. Therefore, a direct comparison of our results with experiments is currently not feasible. To the best of our knowledge, the only two homometallic rings containing eight V magnetic centers ever synthesized and characterized experimentally are reported in Refs. [61] and [62]. However, compared to our V8, they feature a different organic scaffold around the TM centers, and, quite interestingly, report different ground-state magnetic configurations, FM and AFM, respectively. On the other hand, two heterometallic V-based rings with the same structure as our V8 (see Fig. 2) have been synthesized, with single Ni and Zn impurities (V7Ni and V7Zn, respectively). Their experimental characterization shows an AFM ground state [65, 40, 67] which has led authors to expect a similar behavior also for V8. A comparison of these systems with the ones studied here is left for future work.
@@ III.2 Exchange interactions
The magnetic couplings are evaluated from ab initio calculations, by fitting the DFT total energy of suitably chosen spin configurations with the predictions of the spin Hamiltonian. Besides widely used local and semilocal approximations to the exchange-correlation (xc) functional such as LDA [63, 64] and GGA [68], we use the popular DFT+U [69, 70] and the extended DFT+U+V [71, 72] approaches, necessary to improve the $d$ states localization and achieve a more accurate representation of the magnetic moments and their interactions.
@@
Here we follow a somewhat simpler approach based on the broken-symmetry DFT method (BS-DFT)  [79, 80, 81, 82, 83, 84, 85, 86, 87]. More specifically, we adopt the projected BS-DFT, a method which relies on the same assumptions of the four-states method [76, 77, 78]. This approach assumes that the exchange coupling constants do not depend significantly on the electronic states, whose energy is compared to compute them. Other, more sophisticated methods based on quantum-chemistry, suggest that this might not be true for all systems  [88, 89, 90]. While DFT-based calculations cannot be expected to capture the energetics of multi-configurational states, the Hubbard corrections we employ here should make the energy of single spin-configurations quite reliable. These states are the basis of the BS-DFT approach which consists in mapping DFT total energies of selected spin configurations onto the expectation values of the spin Hamiltonian (Eq. 1) on the same states, i. e. on single-determinants constructed from the appropriate eigenstates of $\hat{S}^{z}$ on each magnetic ion. In practice, if we indicate these states as |$\ket{\sigma_{1}^{\left(k\right)},\dots,\sigma_{8}^{\left(k\right)}}$, where $\sigma_{i}^{\left(k\right)}$ is the eigenvalue of $\hat{S}^{z}$ on the $i^{th}$ magnetic site for the $k^{th}$ spin configuration, our approach corresponds to identifying $E_{DFT}^{(k)}$ with $\bra{\sigma_{1}^{\left(k\right)},\dots,\sigma_{8}^{\left(k\right)}}\hat{H}_{ex}\ket{\sigma_{1}^{\left(k\right)},\dots,\sigma_{8}^{\left(k\right)}}$. When a sufficiently large number of states is considered (typical pools of states include FM, AFM and some intermediate configurations), the effective exchange parameters of $\hat{H}_{ex}$ can be extracted from the best fit of the DFT total energies.
@@
After having computed the magnetic couplings, it is worth stressing that, while the structure of the molecule possesses a D4 symmetry, due to the alternate modulation of ligands groups, the magnetic centers are found to be approximately equivalent and deviations from a D8 symmetry can be safely neglected both for the modulus of the localized spins and for the interactions between them. In addition, due to the annular geometry of the systems considered here, periodic boundary conditions are assumed on the site indexes: $i+8=i$.
@@ III.3 Computational details
Density Functional Theory (DFT) calculations were performed using the publicly available, plane-wave pseudopotential package Quantum ESPRESSO (QE) [91, 92, 93]. Spin-resolved, Hubbard-corrected calculations were based on a collinear-spin formulation of DFT. Specifically, we compare results obtained from two most popular approximations for the exchange-correlation (xc) functional: the local density approximation (LDA) [63, 64] and the PBE flavor of the generalized-gradient approximation (GGA) [68]. The pseudopotentials of the TMs were obtained from the QE PSlibrary archive [94, 95] and were chosen to be Ultrasoft (US) and projected-augmented wave (PAW) for GGA and LDA, respectively. Kohn-Sham wavefunctions and charge density were expanded in plane-waves up to a kinetic-energy cutoffs of 80Ry and 640Ry, respectively, in all cases.
Due to the complexity of the real molecular systems, a simplification of the structures was necessary to make the calculations affordable. This was achieved using the widely adopted “hydrogen saturation” procedure [43, 55, 44], which consists in replacing terminal methyl (-CH3) groups with single H atoms. The procedure was sequentially applied twice, and the resulting structure can be observed in Fig. 5, to be compared with Fig. 2.
Figure 5: Reduced molecular structure, see text for explanation.
@@
In order to improve the $d$ states localization and achieve a better representation of magnetic properties, we also used the DFT+U and DFT+U+V schemes. On-site $U$ and inter-site $V$ Hubbard parameters were computed using linear-response theory [96] within the automatic approach introduced in Refs. [97, 98, 99, 100]. To speed-up this calculation, a linear model of the rings’ local structure was used, i.e. a chain containing only two magnetic centers instead of eight, with the same octahedral environment of the TM centers as in the whole molecule.
Figure 6: Linearized chain model of the ring molecule, used to compute the Hubbard parameters. The vertical lines delimitate the two-magnetic-sites unit cell, see main text.
The chain is obtained from the simplified ring in Fig. 5 by changing the position of the ligand groups around the magnetic centers in such a way they can be arranged in a linear conformation. In the process, the structure of the local O-F octahedra around the TMs is maintained as close as possible to the original one. The main difference between the local structure around the TM centers in the ring and in the chain is in the disposition of the F anions that are not anymore all on the same side towards the internal hole of the ring, but assume alternate positions on either side of the chain axis. Due to this alternate structure the shortest chain that can be used contains two TM units, which is the system used for the calculation of the Hubbard parameters. Test calculations of the Hubbard parameters in the ring and in the chain optimized with GGA, yielded almost coincident results  [^1] A more exhaustive discussion about this procedure and the importance of a comparison between curved and linear molecules, will be given in a separate publication.
In all cases the calculation of these parameters was performed self-consistently with respect to the electronic state: $U$’s and $V$’s were obtained perturbing a DFT+U(+V) ground state until input and output values converged. The obtained values are reported in Table 1, showing, for both systems, the considered xc functionals and flavours of the Hubbard correction (DFT+U and DFT+U+V), the value of the computed Hubbard parameters.
@@
It is important to remark that the results shown in Table 1 refer to the ground-state spin configuration of each system. However, test calculations have shown that the Hubbard parameters’ dependence on the arrangement of magnetic moments can be safely neglected. Table 1 shows relatively minor differences (0.1 - 0.2 eV) in the value of $U$ and $V$ between LDA and GGA, independently of the flavor of the Hubbard correction. In contrast, turning on the inter-site interaction ($V$) leads to a significant upward shift ($\approx$ 1 eV) of the on-site Hubbard $U$. At the same time the intersite $V$ exhibits surprisingly similar values for Cr and V, probably due to the close resemblance of the ligands’ structure around the TMs in the two systems. The on-site Hubbard parameters show instead a more pronounced difference between the two systems with those of Cr $\approx$ 0.6 eV higher than those of V. Finally, LDA yields higher values than GGA; the shift, which is non uniform and slightly more pronounced for $U$ than for $V$, is of the order of 0.1 eV.
@@ IV Results
where the additive constant $c$ fixes the energy reference. The considered spin configurations are identified by the sign of the spin on each magnetic site (the modulus being always maximum: $|\sigma_{i}|=S_{i}$) which is controlled by the proper initialization of the local magnetization. Typically, between the FM and the AFM states we consider five/six intermediate configurations, some exemplified in Fig. 4, featuring various numbers of spin-flips in various positions.
A first rough estimate of the nearest-neighbor couplings can be also obtained from the direct comparison of the FM and AFM spin arrangements, see Fig. 4. The results for LDA and GGA are reported in Table 2, showing
@@
that both functionals yield quite large exchange couplings  [^3] We stress that results in Table 2 are rather sensitive to the molecular structure (they change significantly for minor atomic readjustments) which, in turn, is hard to optimize as the convergence at the electronic level is often difficult, especially for the LDA Cr8 value, not fully converged. This sensitivity disappears when using DFT+U or DFT+U+V (Tables 3 and 4). We thus believe it to be another effect of the overdelocalization of valence $d$ electrons in LDA or GGA.. In the case of Cr8, the values strongly overestimate the experimental one, $\sim$1.46 meV, from inelastic neutron scattering (INS) data [58] and electron paramagnetic resonance (EPR) measurements [42, 41]. The mismatch is larger in LDA, which for V8 yields an unrealistic $\sim$-38 meV value. Besides these quantitative inaccuracies, it is important to notice the difference in the sign of the couplings between V8 and Cr8. In agreement with literature [59, 42, 41, 58, 44, 54, 43, 55, 56] Cr8 is found to be AFM, driven by positive nearest-neighbor exchange interactions $J^{(1)}$. In contrast, a completely different picture emerges for V8, exhibiting a negative $J^{\left(1\right)}$ that stabilizes a FM order.
Importantly, using Hubbard corrections strongly modifies the exchange interactions. In the following we present the values obtained with DFT+U and DFT+U+V for both systems, where DFT is either LDA or GGA. In this case the fit on DFT total energy was based on a larger number of spin configurations (see Fig. 4 for a selection) which also gave access to the exchange couplings between second nearest neighbors. Those between further neighbors were found negligible in all cases.
@@
Table 3 collects the couplings obtained from LDA+U and GGA+U, with on-site corrections only. While the couplings upon the Hubbard correction are hugely renormalized as compared to LDA and GGA, the signs of the nearest neighbor couplings are not affected, meaning that the character of the ground state in both systems remains the same: AFM for Cr8 ($J^{\left(1\right)}$ $>$ 0) and FM for V8 ($J^{\left(1\right)}$ $<$ 0). The values of $J^{\left(1\right)}$ are now of the expected order of magnitude and are much closer to the experimental value for Cr8 ($\sim$1.46 meV [58, 41, 42]). Both LDA+U and GGA+U show somewhat underestimated values.
Moreover, besides a FM $J^{\left(1\right)}$, V8 also features a sizeable second-nearest neighbor coupling $J^{\left(2\right)}$, that is positive and, in magnitude, represents a significant fraction of $J^{\left(1\right)}$. In fact, a sizeable $J^{\left(2\right)}$ is needed in order to fit the DFT total energies and make the system ”Heisenberg-like”, i.e., with couplings independent from the spin configuration. $J^{\left(2\right)}$ is also present for Cr8 but is one order of magnitude smaller than for V8. The presence of a relatively large and positive $J^{\left(2\right)}$ is an interesting feature of V8: competing (opposite sign) magnetic interactions of comparable strength might induce frustration and/or enhance the system’s susceptibility to external fields, enabling easier switching between different magnetic orders, an aspect that could be exploited in some applications.
@@
The comparison with Table 3 results clarifies immediately that the coupling $V$ between the TMs and the anions in their first shell of neighbors, generally shifts the value of couplings upward, with negative ones becoming less negative. This slightly improves the agreement for Cr8 with experiments compared to the DFT+U case. The shift is quantitatively more consistent for $J^{\left(1\right)}$ than $J^{\left(2\right)}$ and slightly more pronounced for LDA than GGA. Aside from these differences, however, the overall picture remains unchanged: positive (AFM) $J^{\left(1\right)}$’s and negligible $J^{\left(2\right)}$ for Cr8; negative (FM) $J^{\left(1\right)}$ and consistent, positive $J^{\left(2\right)}$ for V8.
In spite of their structural and chemical similarity, as Cr and V are adjacent on the periodic table, the two molecules exhibit radically different ground states: Cr8 shows an AFM order mainly stabilized by $J^{\left(1\right)}$, while V8 displays a FM arrangement resulting from the competition between $J^{\left(1\right)}$, which is FM and numerically prevalent, and $J^{\left(2\right)}$, which is AFM.